3. Fin a) P(z < 2.37) b) P(z > -1.18) c) P(-1.18 < z < 2.37)
Answer #1

a)

=P(Z<2.37) =0.9911 (by normal table)

b)

=P(Z>-1.18) =1-P{Z<-1.18) =1-0.119 =0.881 (by normal table)

c)

=P(-1.18<Z<2.37) =P(Z<2.37)-P(Z<-1.18) =0.9911-0.119 =0.8721 (by using normal table)